### v1

# Inequality in a model of capitalist economy

Abstract
We analyze inequality aspects of the agent-based model of capitalist economy named Social Architecture of Capitalism that has been introduced by Ian Wright. The model contemplates two main types of agents, workers and capitalists, which can also be unemployed. Starting from a state where all agents are unemployed and possess the same initial wealth, the system, governed by a few simple rules, quickly self-organizes into two classes. After a transient, the model reproduces the statistics of many relevant macroeconomic quantities of real economies worldwide, notably the Boltzmann-Pareto regimes of the distributions of wealth and income. We perform extensive simulations testing the role of the model parameters (number of agents, total wealth, and salary range) on the resulting distribution of wealth and income, the social distribution of agents, and other stylized facts of the dynamics. Our main finding is that, according to the model, in an economy where total wealth is conserved and with a fixed average wage, the increase in wealth per capita comes with more inequality.
Keywords: Econophysics, Agent-based model, Gini index, Inequality.
I Introduction
The search of a first-principle model capable of elucidating the emergence and robustness of Pareto’s law [1] in various contexts is a central challenge in the field of Econophysics. Such a model would provide a fundamental understanding of the mechanisms that drive these distributions, transcending mere empirical observations to reveal the possible underlying principles.
In his seminal work “The Social Architecture of Capitalism” [2], Ian Wright explores the intricate dynamics that results in the self-organization of a capitalist society with a given distribution of wealth among its members. Using an agent-based model known as Social Architecture (SA) model, Wright provides a theoretical framework aiming to explain the ubiquitous presence of Pareto’s law, which reflects that a small percentage of a population controls a large portion of a particular resource such as wealth or income. Wright’s work offers insights into how individual interactions and institutional structures collectively shape economic distributions. The model effectively reproduces the 80/20 rule  [^1] Very often regarded as the $80/20$ rule, meaning that 20% of the population owns 80% of the resources; the exact value of the ratio varies among countries or societies., and similar relations, which are key aspects of Pareto’s law, but, most remarkably, it naturally generates the two-pattern shape (Gamma distribution for the low wealth/income, power-law for the high wealth/income) of the wealth or income distributions, as originally observed by Pareto.
Wright compares his results with real data, demonstrating that the model captures a significant portion of the dynamics of the capitalist system. In addition to the two regimes of wealth and income distributions, the model can reproduce the statistics of other economic indicators, such as firm size (power law), firm size growth rate (double exponential), firm demises per month (lognormal), and recession duration (exponential). These patterns emerge naturally and robustly from the simple set of rules that define the model. Hence, the SA model lays the foundation for exploring various economic phenomena, starting from a microscopic (or agent-based) dynamics controlled by a few parameters. Several follow-up studies have been published [4, 5, 6, 7, 8], but the SA model itself has not been thoroughly analyzed.
Our present goal is to investigate aspects related to inequality in income and wealth distributions using the SA model. We aim to conduct a comprehensive examination of the role played by the model parameters as well as certain details of Wright’s original model. Specifically, we seek to answer the following questions: How do the distributions of income and wealth across different classes depend on parameters such as system size, total wealth, and average wage? How robust are these results? Are they universal and, if so, to what extent? Or are they sensitive to key parameters, and if so, to what degree?
By analyzing the social and economic structures that contribute to income disparities and wealth concentration in capitalist societies, we aim to identify the key parameters and dynamic rules that give rise to Pareto-like distributions and other significant economic measurements. Ultimately, this paper seeks to contribute to a broader understanding of inequality and concentration phenomena in socio-economic systems.
II Model
The model implemented in this paper is the original SA proposed by Ian Wright [2]. However, we have chosen a slightly different notation and terminology to align with those commonly used in econophysics and agent-based modeling. The society consists of $N$ agents (labeled $i=1,\dots,N$ [^2] An agent is not necessarily an individual, but can represent other economic entities., where each agent $i$ has a positive integer quantity $w_{i}(t)$ representing wealth, which varies over time  [^3] While this quantity does not need to be an integer, we follow the original proposal, as this distinction is not critical to our analysis.. The population size $N$ and the total wealth of the system $W=\sum_{i=1}^{N}w_{i}(t)$ are conserved, so the wealth per capita $\overline{w}=W/N$ is constant. Agents can be in one of three classes: employees (working class), employers (capitalist class), or unemployed. Then, each agent is characterized by an index $e_{i}\neq i$, which identifies their employer: $e_{i}=j$ if agent $j$ is the employer of agent $i$, and $e_{i}=0$ if agent $i$ is unemployed or employer. Therefore, at any time $t$, the state of the entire economy is defined by the set of pairs $S(t)=\left\{\left(w_{i}(t),e_{i}(t)\right):1\leq i\leq N\right\}$. A firm consists of a set of employees and an employer, which is the only owner of the firm.
In addition to the system size $N$ and the wealth per capita $\overline{w}$, there are two other parameters: the minimum and maximum wages, $p_{a}$ and $p_{b}$, that each employee can receive. These four are the only parameters of the model. Random selections (of agents, wealth aliquots, wages, etc.) are made uniformly, unless otherwise specified.
In the initial state, all agents have the same wealth ($w_{i}(0)={W}/N$) and are all unemployed ($e_{i}(0)=0$). At each time step (which we set to be a month as in the original paper), the following six steps are executed $N$ times, to give each agent the opportunity to be active once per month on average.
  1. 1.
    Agent selection – An agent $i$ is randomly selected. Then, the following rules are applied, being 3, 4 and 6 related to wealth exchanges, while rules 2 and 5 are related to changes of status.
  2. 2.
    Hiring – Only if $i$ is unemployed, then:
    1. (a)
      A potential employer $j$ is selected from the set $H$ of all agents except the employees, with probability $P(j)=w_{j}/\sum_{n\in H}w_{n}$.
    2. (b)
      If $w_{j}\geq\overline{p}$ (average of the distribution from which the wage (salary) is drawn), then: the agent $i$ is hired by $j$. Hence $j$ becomes an employer if it was previously unemployed.
  3. 3.
    Expenditure (on goods and services) – A random integer $w$ in the range $\left[0,w_{k}\right]$ is selected for a random agent $k\neq i$. Then, the following transfer to the market value $V$ occurs:
    $w_{k}\to w_{k}-w,\quad V\to V+w.$ (1)
  4. 4.
    Market revenue (from sales of goods and services) – Only if $i$ is not unemployed, the following transfer from the market value occurs:
    $\;\;\;\;\;V\to V-w,$
    $\begin{cases}w_{i}\to w_{i}+w,&\mbox{if $i$ is employer,}\\ w_{e_{i}}\to w_{e_{i}}+w,&\mbox{if $i$ is employee,}\end{cases}$
    where $w$ is a random integer number $w\in\left[0,V\right]$. In all cases, the quantity $w$ is counted as firm revenue $r_{i}$.
  5. 5.
    Firing – Only if i is an employer, then:
    1. (a)
      The number $m$ of employees to be fired is defined according to the formula $m=max(n_{i}-\frac{w_{i}}{\overline{p}},0)$, where $n_{i}$ is the number of employees of agent $i$.
    2. (b)
      A number $m$ of agents from the list of $i$’s employees are randomly selected to be fired.
    3. (c)
      Furthermore, if all the workers are fired, then the employer $i$ becomes an unemployed.
  6. 6.
    Wage payment – Apply only when agent $i$ is an employer. For each agent $j$ employed by $i$, the following transfer occurs
    $w_{i}\to w_{i}-p$,
    $w_{j}\to w_{j}+p$,
    where $p$ is a discrete amount randomly selected from the interval $[p_{a},p_{b}]$, but if $p>w_{i}$, then a new wage $p$ is selected from the interval $[0,w_{i}]$.
The update is asynchronous, that is, the system state is updated each time each rule is applied. At the end of $N$ iterations of these six rules, $t$ is incremented by one unit (i.e., one month). Unless otherwise stated, the values of the parameters used in the simulations of the model are: system size $N=1000$ agents, wealth per captia $\overline{w}=100$, salary range $[p_{a},p_{b}]=[10,90]$. Typically, the maximum simulation time is $t_{max}=1000$ years, and the values of all state variables are recorded at the end of each year.
Figure 1: Complementary cumulative wealth distribution (black symbols) $F(w)$ and fraction of the population in each class (color plot) vs. wealth $w$. Result for $N\simeq 10^{5}$, $\overline{w}=100$ and wage range $[10,90]$. Data accumulated over 1000 years, recorded at the end of each year for each agent.
III Preliminary analysis
The distribution of the population among the three classes of agents, along with the wealth distribution within each group, which emerges from the dynamics of the model, is displayed in Fig. 1. We observe a significant concentration of wealth within the capitalist class; however, it does not exhibit the typical pattern of wealth inequality —where most agents have no wealth— commonly found in binary exchange models from the existing literature when there is no intervention (such as taxes or assistance programs) [11, 12, 13, 14].
Figure 2: Distribution of wealth $f(w)$, (a) separated by agents’ type in a system of size $N=3200$, and (b) within the class of capitalists and (c) others, for two different population sizes ($N=100$ and $N=3200$). In all cases, $\overline{w}=100$, then for $N=100$ (3200), the total wealth in the system is $10^{4}$ ($3.2\times 10^{5}$). In the inset of (b), we show a detail of the tails in log-log scale. The segment with slope -2 is drawn for comparison. Data were accumulated over 100 realizations of 1000 years.
The cumulative distribution represented in Fig. 1 can be seen from a different perspective, as given by the probability density functions displayed in Fig. 2(a), which shows that, as an outcome of the model, employees and employers reside in separate wealth ranges. In Fig. 2(b), we focus on the distribution within the capitalist class, for two different system sizes. In both cases, the left-skewed shape is qualitatively similar, as well as the tails (shown in the inset). However, we observe a shift in the mode of the distribution to larger values when the number of agents increases, despite the fact that wealth per capita remains the same ($\overline{w}=100$). For both sizes of the system, the mode is well above $\overline{w}$, indicating wealth concentration. Moreover, increasing system size, which is accompanied by a greater injection of wealth into the system, enhances wealth concentration. Concomitantly, the distribution of wealth within the employee’s class (shown in Fig. 2(c)) suffers opposite impact, such that the mode decreases, and the peak narrows, meaning that wealth becomes more homogeneously distributed in the employee’s class.
In the following sections we will take a closer look on this impact of system size and wealth per capita on the economy dynamics. We will also see that a key parameter of the model is $R\equiv\overline{w}/\overline{p}$. In fact, first note that, according to the rules, the dynamics will not be impacted if wealth is scaled by an arbitrary quantity, for example, by $\overline{p}$, which only alters the monetary unit. In most cases, we will set $\overline{p}=100$. Given the chosen initial state, where agents start as unemployed with the same wealth $\overline{w}$, then, $R\geq 1$ is necessary to comply with $w_{j}>\overline{p}$ (rule 2(b)), to start hiring, and to prevent the system from becoming trapped in its initial state.
The remainder of the paper is organized as follows: In Sec. IV, we investigate the wealth distribution in more detail, focusing systematically on the effects of system size $N$, wealth per capita $\overline{w}$ and wage interval. In Sec. V, we conduct a similar analysis for the income distribution. Finally, in Sec. VI, we present a discussion and the main conclusions.
Figure 3: Wealth per capita within each class vs (a) $N$, for $\overline{w}=100$, and (b) $\overline{w}$, for $N=1000$ (b). For each abscissa value, we plot the average (symbols) and standard deviation (bars) computed over 100 realizations of 1000 years. In all cases, the wage range is [10,90]. The missing points for unemployed in (b) is due to the disappearance of this class. The dashed lines are a guide to the eye.
IV Wealth distribution
In this section, we analyze the wealth distribution by varying key model parameters: system size, total wealth, and average wage.
IV.1 Influence of system size $N$
Figure 4: (a) Complementary cumulative distribution of wealth $F(w)$ of the full population and (b) corresponding Lorenz curves, for different system sizes $N$ indicated in the legend, with $\overline{w}=100$ and wage range $[10,90]$. In (a), the line with slope -1 is drawn for comparison. In (b), the inset shows the average annual (over 1000 years) Gini coefficient vs $N$.
In Fig. 3(a), we plot the wealth per capita across different classes as a function of $N$. The overall average wealth (indicated by symbols) shows a weak dependence on $N$. However, the average wealth of capitalists and the unemployed steadily increases with $N$, while that of employees decreases slightly. This trend suggests that as the population grows, capitalists accumulate more wealth and workers become more impoverished, although these effects stabilize at nearly constant levels for large systems. The concurrent increase in average wealth among the unemployed may be linked to firm crises.
Examining the square root of the variance of the wealth distribution (shown by error bars in Fig. 3), we find that it decreases with $N$ for employed and unemployed agents, indicating increased homogeneity within these groups. Conversely, the variance among capitalists rises with $N$, likely due to the extended asymptotic tail seen in Fig. 2, which suggests a divergent second moment. As $N$ increases keeping $\overline{w}$ constant (thus raising total wealth $W$), the range of capitalist wealth broadens, resulting in greater heterogeneity and an increased likelihood of encountering extremely wealthy agents, as illustrated in Fig. 4(a). Notice the change in the asymptotic behavior of $F(w)$ as $N$ increases, where the tail decays more slowly, approaching the law $F(w)\sim 1/w$, indicating a greater concentration of large fortunes within an oligarchic group. This trend is reflected in the Gini index, which increases gradually with $N$ and tends to saturation, as shown in the inset of Fig. 4(b).
IV.2 Influence of the wealth per capita $\overline{w}$
Figure 3(b) shows how the distribution of wealth across the different classes (scaled by $\overline{w}$) is affected when changing the wealth per capita $\overline{w}$. The increase of $\overline{w}$ has a strong effect on the distribution of wealth, notability enhancing the capitalists’ wealth while reducing the wealth in the working and unemployed classes.
Figure 5: (a) Complementary cumulative distribution of wealth $F(w)$ of the full population and (b) corresponding Lorenz curves, for different levels of the wealth per capita $\overline{w}$, indicated in the legend. The population size is $N=1000$ and the wage range $[10,90]$. In (b), the inset shows the average annual Gini coefficient vs $\overline{w}$.
Upon closer examination of the cumulative distribution in Fig. 5(a), it is clear that the tail decays more slowly as $\overline{w}$ increases. Recalling that this second regime is associated with the capitalist class, according to Figs. 1 and 2, the observed flattening implies that capitalist wealth tends to concentrate in a more exclusive oligarchy with the higher $\overline{w}$. In other words, the probability of finding agents with very high wealth increases. This is an expected effect, also captured by the Lorenz curves in Fig. 5(b), showing that an increase in total wealth leads to a noticeable increase of the Gini coefficient, i.e., the inequality increases. This illustrates the common mistake of using a high gross domestic product as an indicator of progress and equality of opportunities!
For sufficiently large values of $\overline{w}$, (e.g., for $\overline{w}\gtrsim 300$ in the figure), there is a drastic change in the profile of $F(w)$, in which the second regime flattens. This implies the appearance of a disconnected very wealthy class. Moreover, let us remark that, in the extreme of $R\equiv\overline{w}/\overline{p}\gg 1$, the system can freeze with the condensation of the workforce in an oligarchy that concentrates most part of the wealth of the system and their firms cannot bankrupt, which is what happens in the case of the largest value of $\overline{w}$ plotted.
IV.3 Influence of the wage interval
Figure 6: Distributions of (a) wealth $F(w)$ and (b) income $F(y)$, for different wage ranges $[p_{a},p_{b}]$ indicated in the legend. The bimodal case corresponds to $p=10$ or $p=90$, with equal probability 1/2. $\overline{w}=100$ (blue), 600 (red), and $N=10^{3}$. Note that, in all cases, $\overline{p}=50$, and the results coincide for same $\overline{w}$.
We analyze various wage ranges, $[p_{a},p_{b}]$, where salaries are uniformly distributed, and explore their impact on the distribution of wealth. By varying the wage intervals while keeping the mean value $\overline{p}$ fixed, we found that the results remain notably invariant to the wage range, as shown in Fig. 6(a). The same holds for bimodal distributions, as illustrated in the figure for $p=10$ and 90, with probability 1/2. We also explored non-uniform salary distributions, finding again no significant differences. Consequently, a significant conclusion not addressed in the original model is that the effect of wages is determined by their mean value $\overline{p}$. Furthermore, the same results are consistently observed for the same ratio $R=\overline{w}/\overline{p}$. Thus, decreasing $\overline{p}$ exerts an identical effect as increasing $\overline{w}$, both enhancing inequality.
V Income distribution
In this section, we focus on the distribution of incomes (earnings), applying a similar analysis as for the wealth distribution. The results reveal similar trends in relation to inequality.
Figure 7: Annual income distribution due to workers’ wages and due to the revenue of firms. We used $N=102400$, $\overline{w}=100$, and salary range $[10,90]$. The segment with slope -1 is drawn for comparison.
We consider income to be the sum of all remunerations received by the agent over a one-year period. A worker receives a wage (rule 6), while a capitalist receives a revenue, directly or through its employees (rule 4). Additionally, if an agent changes status within a year, they can receive both types of earnings: wages and revenues.
The separate distributions of worker’s wages and of firms’ revenues are shown in Fig. 7, exhibiting a long tail for revenues. Therefore, the distribution of total income, in Figs. 8 and 9, has two regimes that can be associated to each source of income.
Figure 8: (a) Annual income distribution $F(y)$ of the full population and (b) corresponding Lorenz curves, for different system sizes $N$ indicated in the legend, with $\overline{w}=100$ and wage range $[10,90]$. In (a), the line with slope -1 is drawn for comparison. In (b), the inset shows the average annual Gini coefficient vs $N$.
V.1 Influence of the system size $N$
In Fig. 8 we can see how the distribution of wealth is affected by the size of the system. In panel (a), we can see how the two-regime distribution evolves as we increase the number of agents, with a Pareto regime similar to that of the wealth distribution. In the inset of panel (b), we observe a slow increase of the inequality with $N$, as measured by the Gini coefficient, similarly to the behavior of wealth distribution with $N$.
V.2 Influence of the wealth per capita $\overline{w}$
Fig. 9 shows how the system responds when we increase the wealth per capita. Similarly to the behavior found in the previous section for the wealth distribution, a strong increase of income inequality also occurs, as can be measured by the Gini coefficient, shown in the inset of panel (b), which originates from the cumulative wealth distribution and Lorenz curves displayed in Figs. 9 (a) and (b), respectively.
Figure 9: (a) Complementary cumulative distribution of annual income $F(y)$, and (b) corresponding Lorenz curves, for different values of the wealth per capita. Gini index vs. $\overline{w}$ is in the inset. The system size is $N=1000$ and the wage range is [10,90].
V.3 Influence of wage interval
Regarding the wage interval, we confirm here the same behavior observed in wealth distributions, which is dependent only on the average wage $\overline{p}$, as shown in Fig. 6(b). Therefore, increasing $\overline{p}$ produces a rapid decay of income inequality.
VI Conclusions and final remarks
We have conducted a detailed study of the so-called Social Architecture (SA) model, introduced by Wright in 2005 [2]. The model is based on a few simple rules that govern status changes and wealth exchanges. The former occur through hiring and firing, while wealth exchange happens in two ways: directly, as employers pay salaries, and indirectly, as (all) agents contribute to a common fund from which employers retrieve their profits.
Wright presented the model with four parameters: system size $N$, wealth per capita $\overline{w}$, and the lower/upper limits of the salary range, $p_{a}$ and $p_{b}$. Regarding system size, we have shown that all relevant results of the model are independent of $N$, even for systems as small as $N=100$. As for the salary range, the results are sensitive only to the average value $\overline{p}=(p_{a}+p_{b})/2$.
Moreover, since the quantities in the rules can be scaled—indicating a choice of monetary unit without altering the dynamics—we conclude that the only remaining relevant parameter is the ratio $R=\overline{w}/\overline{p}$. On the one hand, $R\geq 1$ is required for the dynamics to evolve from the initial state, as agents would otherwise not have enough wealth to hire employees on average. On the other hand, if $R$ is too large, it leads to an anomalous concentration of wealth and a freezing of the dynamics. Remarkably, within the regime $R\geq 1$, this single parameter model yields a two-regime distribution of wealth and income: Boltzmann-Gibbs-like for low values and Pareto-like at the tails.
Figure 10: Uninterrupted duration (in months) of agents in each class vs. $\overline{w}$, in 100 simulations up to $t=100$ years (after a transient of 50 years), for $N=1000$ and wage range [10,90]. Symbols represent the average and error bars the standard deviation.
According to the model’s predictions, increasing total wealth in the system, while keeping wages fixed, does not alleviate inequality. On the contrary, it amplifies wealth concentration within the capitalist class and reduces the average wealth of employees, thereby increasing inequality, as evidenced by the wealth distribution and the Gini index. Similar effects are observed with respect to income. These outcomes align with the argument that unregulated capitalism drives increasing inequality [15]. As we have seen, the only parameter that matters is $R$, therefore, increasing the average wage has the same effect as decreasing the total wealth.
The model yields a high level of mobility between classes, as shown in Fig. 10, where we plot the average uninterrupted time each agent spends in a given class as a function of wealth per capita. For $\overline{w}\simeq 100$, the time spent in the worker and capitalist classes is shorter than one year and increases with $\overline{w}$ becoming larger for workers, while the agents remain less time unemployed on average as $\overline{w}$ increases. This mobility may seem unrealistic at first sight. However, note, for instance, that according to data from the Bureau of Labor Statistics of US, 50% of the private establishments fail in the first five years [16]. According to the model, this can be understood by observing the higher concentration of wealth that makes capitalist agents more robust against decaying to the unemployed class. Let us also note the homogeneity of the agents, in the sense that they are statistically equivalent and will visit equally the different classes along time, except for very large wealth per capita, in which case freezing occurs.
Some assumptions of the model may be revised, leading to future research perspectives. One such assumption is the expenditure rule. In its present form, an agent could potentially spend all of their wealth in a single transaction. This is clearly extreme when considering a real economy. Modifying this rule could help extend the persistence of capitalists within their class. In fact, this issue is crucial for employers, as the model makes no distinction between personal and firm wealth, meaning that a business owner could exhaust both in a single transaction! In addition, employers hire new employees whenever the opportunity arises, without limitations. It could be interesting to impose constraints related to market value, such as limiting spending to employee savings or firm cash reserves.
Another point is that, in this model, agents are homogeneous and indistinguishable apart from the class to which they belong, with no variation in skill levels. Specifically, hiring assumes that unemployed agents are actively seeking jobs, while employed agents are not pursuing market reallocation, even though people may seek salary upgrades. Some variants addressing this aspect have already been explored [4].
Overall, considering that this is a one-parameter model (governed by $R$), the results are remarkable, offering valuable perspectives on the dynamics of wealth distribution and inequality within a simplified economic framework.
Acknowledgments
We acknowledge financial support from from CAPES, grant #88887.969088/2024-00 and finance code 001. C.A. acknowledges partial financial support from Brazilian agencies CNPq (311435/2020-3) and FAPERJ (CNE E-26/201.109/2021). S.G. would also like to thank CNPq for partially supporting this work under grant #314738/2021–5.